OpenStudy (anonymous):

If 3 copier machines can copy 300 sheets in 3 hours, assuming the same rate, how long (in hours) will it take 6 such copiers to copy 600 sheets?

OpenStudy (anonymous):

3 hours

OpenStudy (anonymous):

3 hours?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

100 pages each. 100*6=600

OpenStudy (anonymous):

300/3=100 pages printed per hour each printer

OpenStudy (anonymous):

retriceuming the same rate! ho ho ho

OpenStudy (anonymous):

Each copier prints 33 sheets each hour. So 6 will produce 6*33 each hour. So 600/(6*33) = 100/33 which equals 3 hours.

OpenStudy (anonymous):

is that some sort of censoring lol.